OpenStudy (anonymous):

The height of a cylinder with a fixed radius of 4 cm is increasing at the rate of 2 cm/min. Find the rate of change of the volume of the cylinder (with respect to time) when the height is 14cm.

OpenStudy (compassionate):

@dan815

OpenStudy (xapproachesinfinity):

well you v=pir^2h take d/dt(v) and plug in your numbers

OpenStudy (xapproachesinfinity):

\(\Huge\frac{d}{dt}(V)=\pi r^2\frac{d}{dt}(h)\) r is fixed r=4

OpenStudy (anonymous):

@xapproachesinginity What do you do after you have got 16pi d/dt(h). Or is 16pi your answer

OpenStudy (xapproachesinfinity):

no you have h' as well it is given \(\large \frac{d}{dt}(h)=2 ~cm/min\)

OpenStudy (xapproachesinfinity):

they said the height is increasing at a rate of 2 cm/min

OpenStudy (anonymous):

so it would be 8pi?

OpenStudy (xapproachesinfinity):

why?m 2*16pi

OpenStudy (xapproachesinfinity):

do you got it?

OpenStudy (anonymous):

so the answer would be 32pi

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

@JFraser can u help me?

OpenStudy (anonymous):

@phi @paki @Zarkon @Destinymasha @agent0smith

OpenStudy (jfraser):

I don't remember this kind of calculus well enough to help, sorry